OpenStudy (anonymous):

Help please! Write explicitly the subgroup of G generated by g: G= S6 and g=(123)(46)

OpenStudy (jamesj):

Well, figure out what is g^2, and g^3, g^4, g^5. We know that g^6 = 1, the identity, so there's only a few elements here

OpenStudy (jamesj):

e.g., g^2 = (123)(46)(123)(46) = (123)(123)(46)(46) Now (46)^2 = 1 (because 4 --> 6 --> 4 and 6 --> 4 --> 6) What then is (123)^2 ?

OpenStudy (anonymous):

is it (132)?

OpenStudy (jamesj):

yes, exactly. So g^2 = (132). Now what is g^3?

OpenStudy (anonymous):

so is the subgroup just {g, g^2...g^6}? whatever the values are?

OpenStudy (jamesj):

(For the record, I'm going to change notation. Using 1 for the identity here is confusing; so let's use the standard e instead.) Yes...

OpenStudy (jamesj):

Given a group G and a member of that group g, then the subgroup generated by g is { g^n | n in Z } In this case, we know that g^6 = e and therefore this set is at most six members {e, g, g^2, g^3, g^4, g^5}

OpenStudy (jamesj):

because g^n = g^(6+n) for every integer n, whether positive or negative.

OpenStudy (jamesj):

and we know that because we are in S6

OpenStudy (jamesj):

Hence it's sufficient to calculate the values of g, g^2, g^3, g^4, g^5 and see whether or not they are unique.

OpenStudy (jamesj):

So that's why we're stepping through these calculations now to see what these group members are

OpenStudy (anonymous):

ok. why is it important that they're unique?

OpenStudy (jamesj):

Only because when we come to write down the set which represents the sub-group, we don't list any element twice.

OpenStudy (jamesj):

For example, consider the sub-group generated by h = (123). Well h^2 = (132) and h^3 = e.

OpenStudy (jamesj):

We can keep on calculating if we want to get h^4 and h^5. But as h^3 = e, it follows that h^4 = h and h^5 = h^2 and there we don't need to do the calculations.

OpenStudy (jamesj):

Therefore the subgroup generated by h = (123) is <(123)> = { e, (123), (132) }

OpenStudy (anonymous):

Great. Thank you!

OpenStudy (jamesj):

Correction/Clarification: it's not true that for any g in S6 that g^6 = e. Although it is true for your particular g. The order of the group Sn -- the permutations of {1, 2, ..., n} -- is n!